# UltraBones100k: A reliable automated labeling method and large-scale dataset for ultrasound-based bone surface extraction

### v2

@@ Abstract
Background: Ultrasound-based bone surface segmentation is crucial in computer-assisted orthopedic surgery. However, ultrasound images have limitations, including a low signal-to-noise ratio, acoustic shadowing, and speckle noise, which make interpretation difficult. Existing deep learning models for bone segmentation rely primarily on costly manual labeling by experts, limiting dataset size and model generalizability. Additionally, the complexity of ultrasound physics and acoustic shadow makes the images difficult for humans to interpret, leading to incomplete labels in low-intensity and anechoic regions and limiting model performance. To advance the state of the art in ultrasound bone segmentation and establish effective model benchmarks, larger and higher-quality datasets are needed.
Methods: We propose a methodology for collecting ex-vivo ultrasound datasets with automatically generated bone labels, including anechoic regions. The proposed labels are derived by accurately superimposing tracked bone Computed Tomography (CT) models onto the tracked ultrasound images. These initial labels are refined to account for ultrasound physics. To clinically evaluate the proposed method, an expert physician from our university hospital specialized on orthopedic sonography assessed the quality of the generated bone labels. A neural network for bone segmentation is trained on the collected dataset and its predictions are compared to expert manual labels, evaluating accuracy, completeness, and F1-score.
Results: We collected the largest known dataset of 100k ultrasound images of human lower limbs with bone labels, called UltraBones100k. A Wilcoxon signed-rank test with Bonferroni correction confirmed that the bone alignment after our optimization pipeline significantly improved the quality of bone labeling ($p<0.001$). The model trained on UltraBones100k consistently outperforms manual labeling in all metrics, particularly in low-intensity regions (320% improvement in completeness at a distance threshold of 0.5 mm)
Conclusion: This work is promising to facilitate research and clinical translation of ultrasound imaging in computer-assisted interventions.
@@ 1 Introduction
In this work, we propose a methodology that enables the collection of large-scale ultrasound datasets with automatically generated bone labels. Instead of relying on manual labeling, our approach utilizes optical tracking to align a 3D bone model, extracted from a preoperative CT scan, with tracked ultrasound images. The initial alignment is further refined using our proposed intensity-based optimization strategy for each frame, rather than relying on traditional global optimization [23]. This approach enhances the precision and completeness of bone labels in ultrasound images. With this methodology, we collected a dataset of 100k annotated ultrasound images with bone labels from 14 human cadaveric lower legs. A deep learning model for bone segmentation is trained on the resulting UltraBones100k dataset, which significantly outperforms manual segmentation performed by an expert surgeon.
To ensure reproducibility of the results, the dataset and pretrained model weights, have been made publicly available [^1] Link will be available upon acceptance..
@@ 2 Related Work
Ultrasound Bone Segmentation. Numerous intensity-based bone segmentation methods have been developed [21, 22]. However, intensity-based methods face challenges in terms of robustness and accuracy. To mitigate these issues, frequency-based methods have been introduced, including phase symmetry and structured phase symmetry techniques [15, 14], which aim to enhance bone boundaries. With the advent of deep learning, neural network models such as CNNs, U-nets, and GANs have been employed for bone segmentation in ultrasound images, targeting various anatomical regions such as the spine [16, 17, 7], pelvis [18, 8], femur [19, 8, 20], tibia [8, 20, 19], fibula [19], and radius [20, 19]. These deep learning methods have demonstrated superior performance over traditional approaches in terms of accuracy, robustness, and runtime. For example, average bone surface distance errors have been reported to range from 0.2 mm to 1.7 mm [20, 24, 7, 18]. Based on accurate 2D bone segmentation results, the reconstructed bone surfaces achieve sub-millimetric one-sided CD [11]. However, these studies rely on manually annotated bone labels from expert surgeons. This process is time-consuming, costly, and requires specialized knowledge of ultrasound imaging and bone anatomy. In addition, different surgeons may also provide different labels for the same ultrasound image, leading to inconsistencies. As a result, the quality and the size of training datasets is limited, typically ranging from dozens to a few thousands of labeled images [22]. These constraints reduce model performance and generalizability. In addition, the absence of standardized evaluation metrics, with over 18 different metrics in use [21], further hinders effective model benchmarking and comparison.
@@ 3.1 Dataset Collection and Automatic Bone labeling
Acquisition Setup. Our dataset acquisition setup using human cadaveric specimens is illustrated in Figure 1. Ethics approval has been granted for our study by the Zurich cantonal ethics committee under BASEC Nr. 2023-01681. Prior to data acquisition, K-wires with a diameter of 2.5 mm and a length of 150 mm (DePuy Synthes, USA), along with custom-made 3D-printed marker mounts equipped with infrared-sensitive fiducial markers, were mounted to each specimen. The K-wires were used to stabilize the bones of interest, including the lower leg and foot, to prevent bone movement during scanning. Optical markers were tracked by an optical tracking system (FusionTrack 500, Atracsys, Switzerland). A CT scan was performed on each specimen with an image resolution of 512 × 512 pixels, an in-plane resolution of 0.839 mm × 0.839 mm, and a slice thickness of 0.6 mm (NAEOTOM Alpha, Siemens, Germany). Bone segmentation was generated using the region growing and thresholding functionalities in Mimics (Materialise, Leuven, Belgium). For data acquisition, each specimen was scanned nine times using a tracked ultrasound device (AixplorerUltimate, SuperSonic Imagine, France) equipped with a linear probe that has an imaging frequency range of 4.0 to 15.0 MHz. The scanning procedure consisted of three forward and backward passes along the longitudinal axis of the tibia, three along the fibula, and three covering the foot (including the bones of the ankle joint such as the calcaneus and talus), aiming to maximize bone surface coverage. Ultrasound images, along with the tracking data of the ultrasound probe and the specimen, were recorded with corresponding timestamps.
@@
    $\mathsf{T}_{\text{US}\leftarrow\text{IP}}=\begin{bmatrix}r_{11}&r_{12}&r_{13}&d_{x}\\ r_{21}&r_{22}&r_{23}&d_{y}\\ r_{31}&r_{32}&r_{33}&d_{z}\\ 0&0&0&1\end{bmatrix}$, belonging to the the 3D Special Euclidean group $\text{SE}(3)$, is the transformation matrix that maps 3D points from the coordinate system of the image plane (origin: top-left corner) to the coordinate system of the marker attached to the ultrasound probe. The determination of this matrix refers to the process of spatial ultrasound calibration [30, 29]. Note that a transformation matrix $\mathsf{T}(\bm{\theta},\mathbf{d})$ can be defined using the Euler rotation angles $\bm{\theta}$ and the translation vector $\mathbf{d}$, which together specify the relationship between two coordinate systems.
@@
Initial Bone Labeling Points are sampled uniformly over the bone CT model to generate the CT point cloud $\mathcal{C}$. Both $\mathcal{C}$ and $\mathcal{U}^{t}$ (with original intensity values) are rendered in Figure 3. The CT-ultrasound bone surface intersection is defined as $\mathcal{I}^{t}=\mathcal{C}\cap\mathcal{U}^{t}=\left\{\mathbf{u}\in\mathcal{U}^{t}\mid\min_{\mathbf{c}\in\mathcal{C}}\lVert\mathbf{u}-\mathbf{c}\rVert\leq\gamma\right\}$ where $\gamma$ is the parameter defining the thickness of the intersection ($\gamma=$0.3 mm determined empirically). Ideally, $\mathcal{I}^{t}$ can be projected back onto the ultrasound image, forming a complete bone label. However, as also reported previously [12], mismatches (1 to 2 mm) often occur due to accumulated errors, including tracking error and calibration error (see Figure 3). Although these mismatches are typically below 2 mm, they are noticeable given the high resolution of ultrasound images, which can reach up to 0.05 mm per pixel.
@@
Optimization Step for Refined Alignment. Bone contours that are roughly perpendicular to the ultrasound probe typically appear as high-intensity regions in ultrasound images due to the underlying ultrasound physics. Based on this principle, we propose an intensity-based optimization method designed to refine the location of CT-derived labels. The proposed optimization relies on the following assumptions: (i) the main error sources are tracking and calibration errors, as the only undetermined variables in Eq. (1) are the tracking matrices ($\mathsf{T}^{t+\Delta t}_{\text{CT}\leftarrow\text{OT}}$ and $\mathsf{T}^{t+\Delta t}_{\text{OT}\leftarrow\text{US}}$) and the calibration matrix ($\mathsf{T}_{\text{US}\leftarrow\text{IP}}$); (ii) pixels with high intensity in the vicinity of the non-corrected labels mostly correspond to the true bone contour. Beneath the bone shadow, there are almost uniformly dark regions, as bones possess the highest acoustic impedance compared to the surrounding soft tissues, causing most acoustic waves to be reflected back; (iii) the CT-ultrasound alignment error does not exceed an mean distance of 2 mm, given that the optical tracking system offers sub-millimeter accuracy [^3] https://atracsys.com/fusiontrack-500/ and the calibration matrix achieves a 3D localization error of 1.07 mm. Furthermore, a random image sampled from each ultrasound sweep to measure mismatch yielded a mean error of $1.13\pm 0.606$ mm.
The key principle of our optimization method is to add small perturbations ($\bm{\theta}_{\epsilon}$ and $\mathbf{d}_{\epsilon}$) to the ultrasound image poses in order to maximize the intensity of pixels at the intersection between the ultrasound image and bone CT model while minimizing deviations from the recorded tracking data. We define the perturbed point cloud as $\mathcal{U}^{t}(\bm{\theta}_{\epsilon},\mathbf{d}_{\epsilon})=\mathsf{T}^{t+\Delta t}_{\text{CT}\leftarrow\text{US}}(\bm{\theta}+\bm{\theta}_{\epsilon},\mathbf{d}+\mathbf{d}_{\epsilon})\cdot\mathsf{T}_{\text{US}\leftarrow\text{IP}}\cdot\mathsf{T}_{\text{S}}\cdot\mathcal{P}^{t}$ and the perturbed intersection as $\mathcal{I}^{t}(\bm{\theta}_{\epsilon},\mathbf{d}_{\epsilon})=\mathcal{C}\cap\mathcal{U}^{t}(\bm{\theta}_{\epsilon},\mathbf{d}_{\epsilon})$. The proposed solution is written as:
@@
Clinical Evaluation. To clinically evaluate the quality of bone annotations following our optimization pipeline, an expert physician specialized on orthopedic sonography assessed the quality of both the initial and optimized bone annotations. For comparison with state-of-the-art, we also employed the traditional ”ultrasound to CT registration” algorithm available in the ImFusion Suite [^4] https://www.imfusion.com/ (ImFusion, Munich, Germany) to compute a global transformation matrix for registering the ultrasound data to the CT volume (the default parameter setting yielded the best results in our experiments), rather than optimizing the alignment for each frame as our method does. In our experiment, one frame was randomly sampled from each ultrasound sweep, resulting in a total of 119 frames. Bone annotations were performed using three alignment methods: the initial alignment, the optimized alignment derived from our method, and the bone alignment obtained after the CT-US registration step. The resulting bone alignment outputs from these three methods were presented to an expert physician specialized on orthopedic sonography in random order. The expert evaluated the quality of each bone alignment using the following criteria:

### v3

@@ Abstract
Background: Ultrasound-based bone surface segmentation is crucial in computer-assisted orthopedic surgery. However, ultrasound images have limitations, including a low signal-to-noise ratio, acoustic shadowing, and speckle noise, which make interpretation difficult. Existing deep learning models for bone segmentation rely primarily on costly manual labeling by experts, limiting dataset size and model generalizability. Additionally, the complexity of ultrasound physics and acoustic shadow makes the images difficult for humans to interpret, leading to incomplete labels in low-intensity and anechoic regions and limiting model performance. To advance the state-of-the-art in ultrasound bone segmentation and establish effective model benchmarks, larger and higher-quality datasets are needed.
Methods: We propose a methodology for collecting ex-vivo ultrasound datasets with automatically generated bone labels, including anechoic regions. The proposed labels are derived by accurately superimposing tracked bone Computed Tomography (CT) models onto the tracked ultrasound images. These initial labels are refined to account for ultrasound physics. To clinically evaluate the proposed method, an expert physician from our university hospital specialized in orthopedic sonography assessed the quality of the generated bone labels. A neural network for bone segmentation is trained on the collected dataset and its predictions are compared to expert manual labels, evaluating accuracy, completeness, and F1-score.
Results: We collected UltraBones100k, the largest known dataset comprising 100k ex-vivo ultrasound images of human lower limbs with bone annotations, specifically targeting the fibula, tibia, and foot bones. A Wilcoxon signed-rank test with Bonferroni correction confirmed that the bone alignment after our optimization pipeline significantly improved the quality of bone labeling ($p<0.001$). The model trained on UltraBones100k consistently outperforms manual labeling in all metrics, particularly in low-intensity regions (at a distance threshold of 0.5 mm: 320% improvement in completeness, 27.4% improvement in accuracy, and 197% improvement in F1 score)
Conclusion: This work is promising to facilitate research and clinical translation of ultrasound imaging in computer-assisted interventions, particularly for applications such as 2D bone segmentation, 3D bone surface reconstruction, and multi-modality bone registration.
@@ 1 Introduction
Bone surface registration between CT and ultrasound has been employed to evaluate the accuracy of bone surface extraction in ultrasound images [6]. However, achieving accurate and fully automatic bone labeling based on CT–ultrasound alignment remains challenging due to the high precision required. In this work, we propose a methodology that enables the collection of large-scale ultrasound datasets with automatically generated bone labels. Instead of relying on manual labeling, our approach utilizes optical tracking to align a 3D bone model, extracted from a preoperative CT scan, with tracked ultrasound images. The initial alignment is further refined using our proposed intensity-based optimization strategy for each frame, rather than relying on traditional global optimization [23]. This approach enhances the precision and completeness of bone labels in ultrasound images. With this methodology, we collected a dataset of 100k annotated ultrasound images with bone labels from 14 human cadaveric lower legs. A deep learning model for bone segmentation is trained on the resulting UltraBones100k dataset, which significantly outperforms manual segmentation performed by an expert surgeon.
To ensure reproducibility, the dataset and pretrained model weights, have been made publicly available at https://github.com/luohwu/UltraBones100k.git.
@@ 2 Related Work
Ultrasound Bone Segmentation. Numerous intensity-based bone segmentation methods have been developed [21, 22]. However, intensity-based methods face challenges in terms of robustness and accuracy. To mitigate these issues, frequency-based methods have been introduced, including phase symmetry and structured phase symmetry techniques [24, 15, 14], which aim to enhance bone boundaries. With the advent of deep learning, neural network models such as CNNs, U-nets, and GANs have been employed for bone segmentation in ultrasound images, targeting various anatomical regions such as the spine [16, 17, 7], pelvis [18, 8], femur [19, 8, 20], tibia [8, 20, 19], fibula [19], and radius [20, 19]. These deep learning methods have demonstrated superior performance over traditional approaches in terms of accuracy, robustness, and runtime. For example, average bone surface distance errors have been reported to range from 0.2 mm to 1.7 mm [20, 25, 7, 18]. Based on accurate 2D bone segmentation results, the reconstructed bone surfaces achieve sub-millimetric one-sided Chamfer Distance (CD) [11]. However, these studies rely on manually annotated bone labels from expert surgeons. This process is time-consuming, costly, and requires specialized knowledge of ultrasound imaging and bone anatomy. In addition, different surgeons may also provide different labels for the same ultrasound image, leading to inconsistencies. As a result, the quality and the size of training datasets is limited, typically ranging from dozens to a few thousands of labeled images [22]. These constraints reduce model performance and generalizability. In addition, the absence of standardized evaluation metrics, with over 18 different metrics in use [21], further hinders effective model benchmarking and comparison.
@@ 3.1 Dataset Collection and Automatic Bone labeling
Cadaveric Specimens. The donated specimens used in this study were fresh-frozen human cadaveric lower limbs, obtained from Science Care (Phoenix, Arizona, USA). The donors cover different physical characteristics, with heights ranging from 155 to 185 cm, weights from 59 to 99 kg, and BMI values between 19 and 30. The age range spans from 41 to 76 years, and the dataset includes 7 male and 7 female specimens. Furthermore, arthritis status is documented, with some specimens having a history of falls or related conditions. Following donor death, the limbs were frozen shortly thereafter and subsequently stored in a dedicated laboratory freezer at $-22\,^{\circ}\mathrm{C}$. Prior to use in the experimental setup, the specimens were thawed at room temperature ($\sim 20\,^{\circ}\mathrm{C}$) for 36 hours, in accordance with an internal standardized wetlab protocol. This thawing process was performed to ensure complete tissue thawing while minimizing tissue degradation and avoiding thermal artifacts that could affect ultrasound imaging. To maintain tissue integrity, all specimens were handled exclusively by wetlab-trained specialist medical professionals. Thus, all ultrasound scanning was conducted on fully thawed, all ultrasound scanning was performed on fully thawed, fresh-frozen human tissue under controlled and standardized conditions, ensuring high fidelity and reproducibility of imaging results.
@@
Acquisition Setup. Our dataset acquisition setup using human cadaveric specimens is illustrated in Figure 1. Ethics approval has been granted for our study by the Zurich cantonal ethics committee under BASEC Nr. 2023-01681. Prior to data acquisition, K-wires with a diameter of 2.5 mm and a length of 150 mm (DePuy Synthes, USA), along with custom-made 3D-printed marker mounts equipped with infrared-sensitive fiducial markers, were mounted to each specimen. The K-wires were used to stabilize the bones of interest, including the lower leg and foot, to prevent bone movement during scanning. Optical markers were tracked by an optical tracking system (FusionTrack 500, Atracsys, Switzerland). A CT scan was performed on each specimen with an image dimension of 512 × 512 pixels, an in-plane pixel spacing of 0.839 mm × 0.839 mm, and a slice thickness of 0.6 mm (NAEOTOM Alpha, Siemens, Germany). CT Bone segmentation was generated using the region growing and thresholding functionalities in Mimics (Materialise, Leuven, Belgium). For data acquisition, each specimen was scanned nine times using an ultrasound device, the Aixplorer Ultimate (SuperSonic Imagine, Aix-en-Provence, France), equipped with an SL18-5 linear probe (SuperSonic Imagine, Aix-en-Provence, France). The probe features 256 elements, an acoustic footprint of 50 mm, and a bandwidth range of 5–18 MHz. Details of the imaging configuration are presented in Table 1. Multiple 3D-printed optical marker arrays are attached on the probe, optimizing the tracking performance in 6 degrees-of-freedom in the target space. In addition, to ensure proper contact between the probe and the specimens for high-quality ultrasound imaging, a gel pad (aiSonTM Technologies, Zurich, Switzerland) [^2] https://www.aisontechnologies.com/aison-focus-standoff-pad/ was attached to the top of the ultrasound probe. The scanning procedure consisted of three forward and backward passes along the longitudinal axis of the tibia, three along the fibula, and three covering the foot (including the bones of the ankle joint such as the calcaneus and talus), aiming to maximize bone surface coverage. During the scanning, the movement of the ultrasound probe is not limited explicitly. The clinician can keep scanning until our system provide a signal indicating that the probe is outside of the workspace visible by the tracking system, which is considered as the limitation along that direction. Ultrasound images, along with the tracking data of the ultrasound probe and the specimen, were recorded with corresponding timestamps.
Gain 48% Tissue Tuner 1540 m/s
Max Depth 5 cm Focus Depth 2.5 cm
Dynamic Range 62 dB Sector Size Large
Harmonic Imaging Disabled Super Compound Enabled
Needle PLUS Disabled Res/Gen/Pen Res
Table 1: Detailed imaging configuration. Initially, the device’s built-in preset ”Orthopedics” was used for setup. Subsequently, the clinician fine-tuned the parameters listed in the table.
@@
    $\mathsf{T}_{\text{US}\leftarrow\text{IP}}=\begin{bmatrix}r_{11}&r_{12}&r_{13}&d_{x}\\ r_{21}&r_{22}&r_{23}&d_{y}\\ r_{31}&r_{32}&r_{33}&d_{z}\\ 0&0&0&1\end{bmatrix}$, belonging to the 3D Special Euclidean group $\text{SE}(3)$, is the transformation matrix that maps 3D points from the coordinate system of the image plane (origin: top-left corner) to the coordinate system of the marker attached to the ultrasound probe. The determination of this matrix refers to the process of spatial ultrasound calibration [31, 30]. Note that a transformation matrix $\mathsf{T}(\bm{\theta},\mathbf{d})$ can be defined using the Euler rotation angles $\bm{\theta}$ and the translation vector $\mathbf{d}$, which together specify the relationship between two coordinate systems.
@@
Initial Bone Labeling Points are sampled uniformly over the bone CT model to generate the CT point cloud $\mathcal{C}$. Both $\mathcal{C}$ and $\mathcal{U}^{t}$ (with original intensity values) are rendered in Figure 3. The CT-ultrasound bone surface intersection is defined as $\mathcal{I}^{t}=\mathcal{C}\cap\mathcal{U}^{t}=\left\{\mathbf{u}\in\mathcal{U}^{t}\mid\min_{\mathbf{c}\in\mathcal{C}}\lVert\mathbf{u}-\mathbf{c}\rVert\leq\gamma\right\}$ where $\gamma$ is the parameter defining the thickness of the intersection ($\gamma=$0.3 mm determined empirically). Ideally, $\mathcal{I}^{t}$ can be projected back onto the ultrasound image, forming a complete bone label. However, as also reported previously [12], mismatches (1 to 2 mm) often occur due to accumulated errors, including tracking error and calibration error (see Figure 3). Although these mismatches are typically below 2 mm, they are noticeable given low pixel spacing of ultrasound images, which can be as small as 0.05 mm per pixel.
@@
Optimization Step for Refined Alignment. Bone contours that are roughly perpendicular to the ultrasound probe typically appear as high-intensity regions in ultrasound images due to the underlying ultrasound physics. Based on this principle, we propose an intensity-based optimization method designed to refine the location of CT-derived labels. The proposed optimization relies on the following assumptions: (i) the main error sources are tracking and calibration errors, as the only undetermined variables in Eq. (1) are the tracking matrices ($\mathsf{T}^{t+\Delta t}_{\text{CT}\leftarrow\text{OT}}$ and $\mathsf{T}^{t+\Delta t}_{\text{OT}\leftarrow\text{US}}$) and the calibration matrix ($\mathsf{T}_{\text{US}\leftarrow\text{IP}}$); (ii) pixels with high intensity in the vicinity of the non-corrected labels mostly correspond to the true bone contour. Beneath the bone shadow, there are almost uniformly dark regions, as bones possess the highest acoustic impedance compared to the surrounding soft tissues, causing most acoustic waves to be reflected back; (iii) the CT-ultrasound alignment error does not exceed a mean distance of 2 mm, given that the optical tracking system offers sub-millimeter accuracy [^3] https://atracsys.com/fusiontrack-500/ and the calibration matrix achieves a 3D localization error of 1.07 mm. Furthermore, a random image sampled from each ultrasound sweep to measure mismatch yielded a mean error of $1.13\pm 0.606$ mm.
The key principle of our optimization method is to add small perturbations ($\bm{\theta}_{\epsilon}$ and $\mathbf{d}_{\epsilon}$) to each ultrasound image pose in order to maximize the intensity of pixels at the intersection between the ultrasound image and bone CT model while minimizing deviations from the recorded tracking data. We define the perturbed point cloud as $\mathcal{U}^{t}(\bm{\theta}_{\epsilon},\mathbf{d}_{\epsilon})=\mathsf{T}^{t+\Delta t}_{\text{CT}\leftarrow\text{US}}(\bm{\theta}+\bm{\theta}_{\epsilon},\mathbf{d}+\mathbf{d}_{\epsilon})\cdot\mathsf{T}_{\text{US}\leftarrow\text{IP}}\cdot\mathsf{T}_{\text{S}}\cdot\mathcal{P}^{t}$ and the perturbed intersection as $\mathcal{I}^{t}(\bm{\theta}_{\epsilon},\mathbf{d}_{\epsilon})=\mathcal{C}\cap\mathcal{U}^{t}(\bm{\theta}_{\epsilon},\mathbf{d}_{\epsilon})$. The proposed solution is written as:
@@
Clinical Evaluation. To clinically evaluate the quality of bone annotations following our optimization pipeline, tow expert physician specialized in orthopedic sonography assessed the quality of both the initial and optimized bone annotations. For comparison with State-of-the-Art (SOTA), we also employed the traditional ”ultrasound to CT registration” algorithm available in the ImFusion Suite [^4] https://www.imfusion.com/ (ImFusion, Munich, Germany) to compute a global transformation matrix for registering the ultrasound data to the CT volume (the default parameter setting yielded the best results in our experiments), rather than optimizing the alignment for each frame as our method does. In our experiment, one frame was randomly sampled from each ultrasound sweep, resulting in a total of 119 frames. Bone annotations were performed using three alignment methods: the initial alignment, the optimized alignment derived from our method, and the bone alignment obtained after the CT-US registration step. The resulting bone alignment outputs from these three methods were presented to an expert physician specialized in orthopedic sonography in random order. The expert evaluated the quality of each bone alignment using the following criteria: